What should be the ratio of concentrations of acetic acid and acetate ions to have a solution with pH 5.74? [pKa for acetic acid is 4.74].


1
Expert's answer
2022-02-28T11:02:33-0500

as it is a buffer solution

ph = pka +log "\frac{salt}{acid}"

5.74= 4.74+ k

k= 1

log salt /acid =1

salt /acid = 10